OpenStudy (anonymous):

Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3.

OpenStudy (anonymous):

this means that you have to plug in the g(x) function into the f(x) function f(g(x))=2(x-7)+5 now simplify f(g(x))=2x-14+5 f(g(x))=2x-9 now plug in -3 for x f(g(-3))=2(-3)-9 simplify f(g(-3))=-6-9 f(g(-3))=-15

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome :)

OpenStudy (zzr0ck3r):

@brangetsbuckets hit the best response button for @rocker128 if you like the answer.